A chemist has a 60 mL beaker of a 40% solution solution. Write an equation for the concentration of the solution after adding x mL of pure water.
Concentration = ________?

Answer by ptaylor(2198)

let y=percent of the solution after the water is added
Amount of the pure solution remains the same before and after water is added
amount of pure solution=0.4*60=24 ml
soooo y(60+x)=24
y=24/(60+x)
Hope this helps-----ptaylor